Derive the expression for the efficiency of a Carnot cycle for an ideal gas. Hence,

obtain its value for a heat engine operating between fixed temperatures 600 K

and 300K.


1
Expert's answer
2021-03-12T07:29:30-0500

Answer

If temperature of source is T and temperature of sink T' then efficiency is given

"\eta=\frac{w}{Q}\\=\frac{Q-Q'}{Q}\\=1-\frac{Q'}{Q}"

In form of temperature

"\eta=1-\frac{T'}{T}"


For given temperatures efficiency

"\eta=1-\frac{300}{600}\\=0.50"